# Radial Basis Function Techniques for Neural Field Models on Surfaces

### v1

@@ 3.2 Quadrature In Flat Domains
The theory of RBF-QF has its roots in finite difference methods, known as RBF-FD. Both quadrature and differentiation are linear functionals – linear maps into $\mathbb{R}$ – and share a common theoretical foundation. Derivative approximation using RBFs was introduced first [33], and the modern RBF-FD method based on local interpolation was discovered independently by several groups shortly after [66, 59, 58, 16]. Due to this history, some authors refer to RBF-QF as a variant of RBF-FD, though it produces quadrature rather than finite difference weights. We begin with the general theory for approximating linear functionals, then specialize to quadrature. This approach clarifies the distinction between RBF-QF and RBF-FD, and highlights when and why a mesh is needed.
@@
$\mathcal{L}f$ $\approx\sum_{E}\begin{bmatrix}\mathcal{L}\bm{\varphi}&\mathcal{L}\bm{\pi}\end{bmatrix}\begin{bmatrix}\bm{c}\\ \bm{d}\end{bmatrix}$ $=\sum_{E}\begin{bmatrix}\mathcal{L}\bm{\varphi}&\mathcal{L}\bm{\pi}\end{bmatrix}\begin{bmatrix}A&\Pi\\ \Pi^{T}&0\end{bmatrix}^{-1}\begin{bmatrix}\bm{f}\\ \bm{0}\end{bmatrix},$
@@
$\bm{w}_{E}=\begin{bmatrix}A&\Pi\\ \Pi^{T}&0\end{bmatrix}^{-1}\begin{bmatrix}\mathcal{L}\bm{\varphi}\\ \mathcal{L}\bm{\pi}\end{bmatrix},\qquad(\bm{w})_{\bm{x}_{i}}=w_{\bm{x}_{i}}=\sum_{E\colon\bm{x}_{i}\in S(E)}(\bm{w}_{E})_{\bm{x}_{i}}.$ (15)
@@
[…]
@@ 3.3 Quadrature on Smooth Closed Manifolds
The method begins with a set of interpolation nodes $\Xi$ and defines two sets of elements. The first is a planar triangulation $\mathcal{E}=\cup_{j=1}^{m}E_{j}$ with nodes exactly equal to $\Xi$. The method is designed for curved (non-planar) domains where the cortical surface $D$ does not coincide with $\mathcal{E}$. It constructs a second set of elements forming a partition $D=\cup_{j=1}^{m}\mathcal{T}_{j}$, where each $\mathcal{T}_{j}$ is a curved triangular patch on $D$, in contrast to the flat triangles $E_{j}$ in $\mathcal{E}$ (See fig. 3.2). The two families share the interpolation nodes, $\Xi\subseteq D\cap\mathcal{E}$, and are paired such that for each $j$ there exists a bijection $\mu_{j}(E_{j})=\mathcal{T}_{j}$. We defer a precise definition of the mappings to a later section, and simply note that once the families $\{E_{j}\}$, $\{\mathcal{T}_{j}\}$, and $\{\mu_{j}\}$ are in place, one can write, without approximation,
[…]
@@
Figure 3.2: A. The surface of a torus ($D_{T}$, transparent light-green) is sampled at some well-distributed points ($\Xi$, black) which are then used for a triangulation ($\{E_{j}\}_{j=1}^{M}$, white flat triangles). Using the method described in Section 3.3, these planar triangles are projected onto the torus to form a partition of surface triangles ($\{\mathcal{T}_{j}\}_{j=1}^{M}$, outlined in yellow). B. Stencil and projected stencil of a flat triangle ($E_{j}$, white, but appears light green because it is behind the surface) from a surface triangulation. Nearby surface points form the stencil ($S(E_{j})$, red) and are projected along lines (blue) emanating from the projection point, onto co-planar points ($\xi$, blue). The projection point is computed from the vertices of the triangle and approximated edge normals (red arrows). C. A detailed diagram relating a flat triangle $E_{j}$ and its associated surface triangle $\mathcal{T}_{j}$. Points $\bm{x}\in\pi_{j}$ on the plane of $E_{j}$ are projected onto $\bm{\xi}\in\mathcal{T}_{j}$ along the lines through the projection point $p_{j}$.
The mappings $\{\mu_{j}\}$ depend on a set of projection points $\{\bm{p}_{j}\}$, as shown in Figure 3.2. For now, we assume these points are given and explain later how to determine them. For each fixed $j$, let $\pi_{j}$ denote the plane in which the triangle $E_{j}$. We define
@@
By construction, $E_{j}$ and $\mathcal{T}_{j}$ intersect at the nodes $\{\bm{x}_{j,1},\bm{x}_{j,2},\bm{x}_{j,3}\}\subset\Xi$ (See fig. 3.2). The mapping $\mu_{j}$ sends each edge of the planar triangle $E_{j}$ connecting $\bm{x}_{j,k}$ and $\bm{x}_{j,l}$ to the corresponding edge of the curved patch $\mathcal{T}_{j}$, preserving endpoints. Each such edge, together with the projection point $\bm{p}_{j}$, lies in a common cutting plane; thus $\bm{p}_{j}$ is the intersection of three cutting planes.
@@
and $\{\bm{x}_{j,1},\ldots,\bm{x}_{j,q_{j}}\}=S(E_{j})$; that is, the quadrature nodes for $E_{j}$ are the pre-images of the stencil points under $\mu_{j}$ (see Section 3.1 for a definition of $S(E_{j})$). Importantly, the nodes $\{\bm{\xi}_{j,r}\}_{r=1}^{q_{j}}$ lie in the plane $\pi_{j}$, but not necessarily within the triangle $E_{j}$ itself, as illustrated in Figure 3.2. The points $\mu_{j}^{-1}(S(E_{j}))\in\pi_{j}$ now form a planar stencil for $E_{j}$, and we use the procedure outlined in Section 3.2. Specifically, we treat each $E_{j}$ as a flat domain and apply RBF-QF locally, using the basis functions evaluated at the stencil nodes and integrating them over $E_{j}$. This yields weights that, when paired with function evaluations on the surface at $\bm{x}_{j,r}=\mu_{j}(\bm{\xi}_{j,r})$, define a quadrature rule that approximates integrals over the curved domain $D$.
@@ 4 Numerical Experiments for Neural Fields
$\bm{x}(\varphi,\theta)$ $=\begin{bmatrix}x(\varphi,\theta)\\ y(\varphi,\theta)\\ z(\varphi,\theta)\end{bmatrix}=\begin{bmatrix}(R+r\cos\theta)\cos\varphi\\ (R+r\cos\theta)\sin\varphi\\ r\sin\theta\end{bmatrix}.$ (17)
@@
$\text{dist}(\bm{x},\bm{y})=\min_{i,j=-1,0,1}\left\|\bm{x}-\bm{y}+\Lambda\left(i\begin{bmatrix}1\\ 0\end{bmatrix}+j\begin{bmatrix}0\\ 1\end{bmatrix}\right)\right\|.$ (18)
@@
$\text{Gauss}(\bm{x},\bm{y};\sigma)$ $=\frac{1}{2\pi\sigma^{2}}\exp\left[-\frac{1}{2\sigma^{2}}\ \text{dist}^{2}(\bm{x},\bm{y})\right].$ (19)
@@
$f[u]$ $=\bigg{(}1+\exp\big{(}-\gamma(u-\vartheta)\big{)}\bigg{)}^{-1}$ (20)
@@ 4.1 Node selection
We briefly describe our selection of quadrature nodes and triangular meshes in each domain. Rather than detailing each construction, we provide simplified description and refer interested readers to the Code availability section (section 6) for implementation details, including quadrature weight generation, node placement, and time integration. For the square domains $D_{1}$ and $D_{2\pi}$ we compare two node families. The first places interior nodes at the vertices of a regular equilateral triangle tiling, with equally spaced nodes along the boundary. We refer to this as the regular grid, since interior stencils are rotated translates of each other and yield identical weights (Figure 4.1E). The second family distributes random interior nodes using a deterministic point-repulsion process, again placing equally spaced nodes along the boundary. An example is shown in Figure 4.1C. In both cases, we use Delaunay triangulation of the quadrature nodes to define the mesh.
@@ 4.2 Quadrature Experiments on the Unit Square
Figure 4.1: RBF-QF on the Unit Square. A. A single test function as in eq. 22). B. Quadrature node locations for $n=2000$ points, colored by their associated weights. Boundary-adjacent nodes have smaller weights due to asymmetric stencils, and five nodes yield negative weights (black circles). Right: a histogram of weights showing a bimodal distribution—one mode near $0.00051/n=$ 0.0005 and a second associated with boundary clustering. C–D. Relative error (C) and log10 relative error (D) of the quadrature rule applied to Gaussian test functions $f_{\bm{x}_{0}}$ centered throughout the domain. E–F. Same as panels C–D, but using a near-regular triangular grid. Interior weights approach $1/n$ and the error approaches machine precision away from the boundary.
To demonstrate more precisely, we introduce a qualitative test in which a localized test function smooths the quadrature operator. Specifically, we define a family of steep Gaussian test functions $f_{\bm{x}_{0}}:D_{1}\to\mathbb{R}$ by 333 The prefactor of $2\pi 10^{2}$ has been used in the numerical experiment, but it has no impact on the relative error presented below, and can be safely ignored
$f_{\bm{x}_{0}}(\bm{x})$ $=2\pi 10^{2}\;\text{Gauss}(\bm{x},\bm{x}_{0};10)$ (22)
where each function is centered at $\bm{x}_{0}\in D_{1}$. These Gaussians, defined in eq. 19, using the doubly-periodic distance eq. 18, integrate to the same constant regardless of their center. The mass is sharply localized near $\bm{x}_{0}$, with rapid decay in all derivatives away from the center. While the periodic distance is not smooth, the resulting test functions are effectively smooth to machine precision. An example is shown in Figure 4.1A.
Figure 4.1B shows the quadrature nodes, colored by their associated weights. We observe that weights along the boundary are consistently smaller. This is partly because boundary nodes appear in fewer stencils, and partly due to mild clustering near the edges—an intentional design to mitigate boundary-related errors such as the Runge phenomenon. Five nodes have negative weights (highlighted with black circles), though each is small in magnitude. The histogram to the right reveals a bimodal distribution: the larger mode is centered near $0.00051/n=$ 0.0005, consistent with the average area per node, while the smaller mode corresponds to the boundary-adjacent nodes. This empirical structure reflects both the local stencil geometry and the rule’s enforcement of exactness for constant functions.
@@
We use localized Gaussian test functions to visualize how quadrature error varies spatially for different node configurations. In Figure 4.1, we demonstrate this on two node sets in the unit square (See Section 4). Figure 4.1C shows a heatmap of $E(\bm{x}_{0})$, the relative signed quadrature error of our Gaussian test functions $f_{\bm{x}_{0}}$. We again use $n=2000$ quadrature nodes, the basic function $r^{3}$, a stencil size of $k=21$, and append third-degree polynomials. The error is relatively small, varies continuously, and is not systematically positive or negative. Most of the error magnitudes are well below the extremal values, highlighting the importance of testing across a variety of test functions.
@@
Any test function can exhibit surprisingly low error due to the random nature of the quadrature nodes. The function $E(\bm{x}_{0})$ is continuous and oscillatory, and since the quadrature rule integrates constants exactly, the average of $E$ over the domain is zero. The continuity of $E(\bm{x}_{0})$ follows from the fact that $Q_{n}f_{\bm{x}_{0}}$ is a continuous function of $\bm{x}_{0}$, and $Qf_{\bm{x}_{0}}$ is constant in $\bm{x}_{0}$. The average value of $E$ is zero because the quadrature weights are exact for constant functions, a fact that can be verified by a short computation. Thus, there must exist closed curves along which $E(\bm{x}_{0})=0$. We visualize this zero-level set by plotting $\log|E(\bm{x}_{0})|$ in Figure 4.1D. While the overall structure remains similar across different meshes, the precise locations of high and low error regions will vary. This implies that, even for a fixed test function, quadrature error exhibits mesh-dependent spatial variation—even when meshes have similar $n$ and point densities. For this reason, it is common in RBF quadrature studies to generate multiple random node sets for each choice of mesh parameters.
In practice, the presence of small negative weights does not prevent convergence. As noted earlier, Figure 4.1B shows five such weights, each small in magnitude, and their influence appears minimal in Figure 4.1C–D. It is often proposed that a quadrature rule is stable provided
@@
i.e., if all weights are non-negative [29]. High-order Newton–Cotes rules, for instance, fail this criterion and are indeed unstable. However, this notion of stability seems overly restrictive for our purposes. A more lenient condition–—one that permits convergence—–is that the difference between the total signed and absolute weights remains bounded as $n\to\infty$ [60, 47]. Indeed, the RBF-QF method has been analyzed under this framework [29], and our experiments confirm that it converges robustly even in the presence of small negative weights.
Lastly, Panels E and F of Figure 4.1 illustrate a case where the quadrature rule achieves spectral accuracy away from the boundary. The test functions and algorithmic parameters remain unchanged, but the nodes are now arranged on a near-regular triangular grid. In this symmetric configuration, the error is extremely small away from the boundary and reaches machine precision near the domain center. The white pixels visible in Panel F result from floating-point coincidence: the analytic integral and quadrature evaluation yield exactly the same double-precision value.
This level of accuracy arises from geometric regularity. When the mesh elements and their associated stencils are rotated translates of each other, the resulting quadrature weights are identical. Coupled with exact integration of constant functions, this symmetry forces all interior weights to equal the average area per node. Importantly, this value is independent of the degree of appended polynomial basis terms, which determine the formal convergence rate. As a result, the quadrature rule achieves arbitrarily high order in the interior—analogous to the spectral accuracy of the trapezoidal rule for periodic functions in one dimension.

### v2

@@ 3.2 Quadrature In Flat Domains
The theory of RBF-QF has its roots in finite difference methods, known as RBF-FD. Both quadrature and differentiation are linear functionals – linear maps into $\mathbb{R}$ – and share a common theoretical foundation. Derivative approximation using RBFs was introduced first [36], and the modern RBF-FD method based on local interpolation was discovered independently by several groups shortly after [69, 62, 60, 19]. Due to this history, some authors refer to RBF-QF as a variant of RBF-FD, though it produces quadrature weights rather than finite difference weights. We begin with the general theory for approximating linear functionals, then specialize to quadrature. This approach clarifies the distinction between RBF-QF and RBF-FD, and highlights when and why a mesh is needed.
@@
$\mathcal{L}f$ $\approx\sum_{E}\begin{bmatrix}\mathcal{L}\bm{\varphi}&\mathcal{L}\bm{\pi}\end{bmatrix}\begin{bmatrix}\bm{c}\\ \bm{d}\end{bmatrix}$ $=\sum_{E}\begin{bmatrix}\mathcal{L}\bm{\varphi}&\mathcal{L}\bm{\pi}\end{bmatrix}\begin{bmatrix}A&\Pi\\ \Pi^{T}&0\end{bmatrix}^{-1}\begin{bmatrix}\bm{f}_{E}\\ \bm{0}\end{bmatrix}$,
@@
(14) $\bm{w}_{E}=\begin{bmatrix}A&\Pi\\ \Pi^{T}&0\end{bmatrix}^{-1}\begin{bmatrix}\mathcal{L}\bm{\varphi}\\ \mathcal{L}\bm{\pi}\end{bmatrix},\qquad(\bm{w})_{\bm{x}_{i}}=w_{\bm{x}_{i}}=\sum_{E\colon\bm{x}_{i}\in S(E)}(\bm{w}_{E})_{\bm{x}_{i}}.$
@@
Since we focus on quadrature, we set $\mathcal{L}f=Qf=\int_{D}f(y)d\mu(y)$ as in Section 2. While our theory suggests $Q_{n}=QP_{n}$, we use the simpler form $Q_{n}=Qs=\bm{w}^{T}\bm{f}$, which offers the same accuracy and is easier to compute. Each functional evaluation $\mathcal{L}\bm{\varphi},\mathcal{L}\bm{\pi}$ thus involves analytically integrating an RBF or polynomial basis function over an element $E$. The resulting quadrature approximation is then $Q_{n}f=\bm{w}^{T}\bm{f}$.
@@ 3.3 Quadrature on Smooth Closed Manifolds
The method begins with a set of interpolation nodes $\Xi$ and defines two sets of elements. The first is a planar triangulation $\mathcal{E}=\cup_{j=1}^{m}E_{j}$ with nodes exactly equal to $\Xi$. The method is designed for curved (non-planar) domains where the cortical surface $D$ does not coincide with $\mathcal{E}$. It constructs a second set of elements forming a partition $D=\cup_{j=1}^{m}\mathcal{T}_{j}$, where each $\mathcal{T}_{j}$ is a curved triangular patch on $D$, in contrast to the flat triangles $E_{j}$ in $\mathcal{E}$ (See Fig. 2). The two families share the interpolation nodes, $\Xi\subseteq D\cap\mathcal{E}$, and are paired such that for each $j$ there exists a bijection $\mu_{j}(E_{j})=\mathcal{T}_{j}$. We defer a precise definition of the mappings to a later section, and simply note that once the families $\{E_{j}\}$, $\{\mathcal{T}_{j}\}$, and $\{\mu_{j}\}$ are in place, one can write, without approximation,
(15) $\int_{D}f(\bm{x})\,dS=\sum_{j=1}^{m}\int_{\mathcal{T}_{j}}f(\bm{x})\,dS=\sum_{j=1}^{m}\int_{E_{j}}f\bigl(\mu_{j}(\bm{\xi})\bigr)\,\|(\partial_{\xi_{1}}\mu_{j}\times\partial_{\xi_{2}}\mu_{j})(\bm{\xi})\|\,d\bm{\xi},$
@@
Figure 2: A. The surface of a torus ($D_{T}$, transparent light-green) is sampled at some well-distributed points ($\Xi$, black) which are then used for a triangulation ($\{E_{j}\}_{j=1}^{M}$, white flat triangles). Using the method described in Section 3.3, these planar triangles are projected onto the torus to form a partition of surface triangles ($\{\mathcal{T}_{j}\}_{j=1}^{M}$, outlined in yellow). B. Stencil and projected stencil of a flat triangle ($E_{j}$, white, but appears light green because it is behind the surface) from a surface triangulation. Nearby surface points form the stencil ($S(E_{j})$, red) and are projected along lines (blue) emanating from the projection point, onto co-planar points ($\xi$, blue). The projection point is computed from the vertices of the triangle and approximated edge normals (red arrows). C. A detailed diagram relating a flat triangle $E_{j}$ and its associated surface triangle $\mathcal{T}_{j}$. Points $\bm{x}\in\pi_{j}$ on the plane of $E_{j}$ are projected onto $\bm{\xi}\in\mathcal{T}_{j}$ along the lines through the projection point $p_{j}$.
The mappings $\{\mu_{j}\}$ depend on a set of projection points $\{\bm{p}_{j}\}$, as shown in Figure 2. For now, we assume these points are given and explain later how to determine them. For each fixed $j$, let $\pi_{j}$ denote the plane in which the triangle $E_{j}$. We define
@@
By construction, $E_{j}$ and $\mathcal{T}_{j}$ intersect at the nodes $\{\bm{x}_{j,1},\bm{x}_{j,2},\bm{x}_{j,3}\}\subset\Xi$ (See Fig. 2). The mapping $\mu_{j}$ sends each edge of the planar triangle $E_{j}$ connecting $\bm{x}_{j,k}$ and $\bm{x}_{j,l}$ to the corresponding edge of the curved patch $\mathcal{T}_{j}$, preserving endpoints. Each such edge, together with the projection point $\bm{p}_{j}$, lies in a common cutting plane; thus $\bm{p}_{j}$ is the intersection of three cutting planes.
@@
and $\{\bm{x}_{j,1},\ldots,\bm{x}_{j,q_{j}}\}=S(E_{j})$; that is, the quadrature nodes for $E_{j}$ are the pre-images of the stencil points under $\mu_{j}$ (see Section 3.1 for a definition of $S(E_{j})$). Importantly, the nodes $\{\bm{\xi}_{j,r}\}_{r=1}^{q_{j}}$ lie in the plane $\pi_{j}$, but not necessarily within the triangle $E_{j}$ itself, as illustrated in Figure 2. The points $\mu_{j}^{-1}(S(E_{j}))\in\pi_{j}$ now form a planar stencil for $E_{j}$, and we use the procedure outlined in Section 3.2. Specifically, we treat each $E_{j}$ as a flat domain and apply RBF-QF locally, using the basis functions evaluated at the stencil nodes and integrating them over $E_{j}$. This yields weights that, when paired with function evaluations on the surface at $\bm{x}_{j,r}=\mu_{j}(\bm{\xi}_{j,r})$, define a quadrature rule that approximates integrals over the curved domain $D$.
@@ 4 Numerical Experiments for Neural Fields
(16) $\bm{x}(\varphi,\theta)$ $=\begin{bmatrix}x(\varphi,\theta)\\ y(\varphi,\theta)\\ z(\varphi,\theta)\end{bmatrix}=\begin{bmatrix}(R+r\cos\theta)\cos\varphi\\ (R+r\cos\theta)\sin\varphi\\ r\sin\theta\end{bmatrix}$.
@@
[…]
@@
(18) $\text{Gauss}(\bm{x},\bm{y};\sigma)$ $=\frac{1}{2\pi\sigma^{2}}\exp\left[-\frac{1}{2\sigma^{2}}\ \text{dist}^{2}(\bm{x},\bm{y})\right].$
@@
(19) $f[u]$ $=\bigg(1+\exp\big(-\gamma(u-\vartheta)\big)\bigg)^{-1}$
@@ 4.1 Node selection
We briefly describe our selection of quadrature nodes and triangular meshes in each domain. Rather than detailing each construction, we provide simplified description and refer interested readers to the Code availability section (Section 6) for implementation details, including quadrature weight generation, node placement, and time integration. For the square domains $D_{1}$ and $D_{2\pi}$ we compare two node families. The first places interior nodes at the vertices of a regular equilateral triangle tiling, with equally spaced nodes along the boundary. We refer to this as the regular grid, since interior stencils are rotated translates of each other and yield identical weights (Figure 3E). The second family distributes random interior nodes using a deterministic point-repulsion process, again placing equally spaced nodes along the boundary. An example is shown in Figure 3C. In both cases, we use Delaunay triangulation of the quadrature nodes to define the mesh.
@@ 4.2 Quadrature Experiments on the Unit Square
Figure 3: RBF-QF on the Unit Square. A. A single test function as in Eq. 21). B. Quadrature node locations for $n=2000$ points, colored by their associated weights. Boundary-adjacent nodes have smaller weights due to asymmetric stencils, and five nodes yield negative weights (black circles). Right: a histogram of weights showing a bimodal distribution—one mode near $1/n=0.0005$ and a second associated with boundary clustering. C–D. Relative error (C) and log10 relative error (D) of the quadrature rule applied to Gaussian test functions $f_{\bm{x}_{0}}$ centered throughout the domain. E–F. Same as panels C–D, but using a near-regular triangular grid. Interior weights approach $1/n$ and the error approaches machine precision away from the boundary.
To demonstrate more precisely, we introduce a qualitative test in which a localized test function smooths the quadrature operator. Specifically, we define a family of steep Gaussian test functions $f_{\bm{x}_{0}}:D_{1}\to\mathbb{R}$ by  [^3] The prefactor of $2\pi 10^{2}$ has been used in the numerical experiment, but it has no impact on the relative error presented below, and can be safely ignored
(21) $f_{\bm{x}_{0}}(\bm{x})$ $=2\pi 10^{2}\;\text{Gauss}(\bm{x},\bm{x}_{0};10)$
where each function is centered at $\bm{x}_{0}\in D_{1}$. These Gaussians, defined in Eq. 18, using the doubly-periodic distance Eq. 17, integrate to the same constant regardless of their center. The mass is sharply localized near $\bm{x}_{0}$, with rapid decay in all derivatives away from the center. While the periodic distance is not smooth, the resulting test functions are effectively smooth to machine precision. An example is shown in Figure 3A.
Figure 3B shows the quadrature nodes, colored by their associated weights. We observe that weights along the boundary are consistently smaller. This is partly because boundary nodes appear in fewer stencils, and partly due to mild clustering near the edges—an intentional design to mitigate boundary-related errors such as the Runge phenomenon. Five nodes have negative weights (highlighted with black circles), though each is small in magnitude. The histogram to the right reveals a bimodal distribution: the larger mode is centered near $1/n=0.0005$, consistent with the average area per node, while the smaller mode corresponds to the boundary-adjacent nodes. This empirical structure reflects both the local stencil geometry and the rule’s enforcement of exactness for constant functions.
@@
We use localized Gaussian test functions to visualize how quadrature error varies spatially for different node configurations. In Figure 3, we demonstrate this on two node sets in the unit square (See Section 4). Figure 3C shows a heatmap of $E(\bm{x}_{0})$, the relative signed quadrature error of our Gaussian test functions $f_{\bm{x}_{0}}$. We again use $n=2000$ quadrature nodes, the basic function $r^{3}$, a stencil size of $k=21$, and append third-degree polynomials. The error is relatively small, varies continuously, and is not systematically positive or negative. Most of the error magnitudes are well below the extremal values, highlighting the importance of testing across a variety of test functions.
@@
Any test function can exhibit surprisingly low error due to the random nature of the quadrature nodes. The function $E(\bm{x}_{0})$ is continuous and oscillatory, and since the quadrature rule integrates constants exactly, the average of $E$ over the domain is zero. The continuity of $E(\bm{x}_{0})$ follows from the fact that $Q_{n}f_{\bm{x}_{0}}$ is a continuous function of $\bm{x}_{0}$, and $Qf_{\bm{x}_{0}}$ is constant in $\bm{x}_{0}$. The average value of $E$ is zero because the quadrature weights are exact for constant functions, a fact that can be verified by a short computation. Thus, there must exist closed curves along which $E(\bm{x}_{0})=0$. We visualize this zero-level set by plotting $\log|E(\bm{x}_{0})|$ in Figure 3D. While the overall structure remains similar across different meshes, the precise locations of high and low error regions will vary. This implies that, even for a fixed test function, quadrature error exhibits mesh-dependent spatial variation—even when meshes have similar $n$ and point densities. For this reason, it is common in RBF quadrature studies to generate multiple random node sets for each choice of mesh parameters.
In practice, the presence of small negative weights does not prevent convergence. As noted earlier, Figure 3B shows five such weights, each small in magnitude, and their influence appears minimal in Figure 3C–D. It is often proposed that a quadrature rule is stable provided
@@
i.e., if all weights are non-negative [31]. High-order Newton–Cotes rules, for instance, fail this criterion and are indeed unstable. However, this notion of stability seems overly restrictive for our purposes. A more lenient condition–—one that permits convergence—–is that the difference between the total signed and absolute weights remains bounded as $n\to\infty$ [64, 50]. Indeed, the RBF-QF method has been analyzed under this framework [31], and our experiments confirm that it converges robustly even in the presence of small negative weights. Alternative strategies, such as $L^{1}$ optimization of stencil weights [1], have been proposed to mitigate or eliminate negative weights while maintaining accuracy, but we do not pursue these modifications here.
Lastly, Panels E and F of Figure 3 illustrate a case where the quadrature rule achieves spectral accuracy away from the boundary. The test functions and algorithmic parameters remain unchanged, but the nodes are now arranged on a near-regular triangular grid. In this symmetric configuration, the error is extremely small away from the boundary and reaches machine precision near the domain center. The white pixels visible in Panel F result from floating-point coincidence: the analytic integral and quadrature evaluation yield exactly the same double-precision value.
This level of accuracy arises from geometric regularity. When the mesh elements and their associated stencils are rotated translates of each other, the resulting quadrature weights are identical. Coupled with exact integration of constant functions, this symmetry forces all interior weights to equal the average area per node. Importantly, this value is independent of the degree of appended polynomial basis terms, which determine the formal convergence rate. As a result, the quadrature rule achieves arbitrarily high order in the interior—analogous to the spectral accuracy of the trapezoidal rule for periodic functions in one dimension. We observe the same phenomenon on other geometries: when nodes are placed on regular, nearly uniform grids on the square (Fig. SM1), sphere (Fig. SM2), or cyclide (Fig. SM3), the quadrature achieves high accuracy, as documented in the Supplementary Materials.